OpenStudy (anonymous):

solve and explain 6(p^2-1)=5p

OpenStudy (joannablackwelder):

First distribute.

OpenStudy (joannablackwelder):

Can you do that?

OpenStudy (anonymous):

\[6p ^{2}-6-5=0?\]

OpenStudy (joannablackwelder):

Close! You left off the p from the 5p. Now do you know how to factor using grouping?

OpenStudy (joannablackwelder):

\[6p ^{2}-5p-6=0\]

OpenStudy (anonymous):

so now I will be able to plug in those numbers into the quadratic formula, correct?

OpenStudy (joannablackwelder):

You could do that. :)

OpenStudy (anonymous):

Thank you!

OpenStudy (joannablackwelder):

No worries. :) Let me know what you get.

OpenStudy (anonymous):

is it X= 3/2 & X= -2/3

OpenStudy (joannablackwelder):

Yup! :)

OpenStudy (anonymous):

Thank you! :)

OpenStudy (joannablackwelder):

No worries.